OpenStudy (anonymous):

If the complex number x = 3 + bi and |x|2 = 13, which is a possible value of b? 2 4 9 10

OpenStudy (anonymous):

|x|^2 = a^2 + b^2 => 13 = 3^2 + b^2 => b^2 = ?

OpenStudy (anonymous):

9

OpenStudy (mathmale):

@elocke: Your " |x|2 = 13 " looks strange and is ambiguous. Does that 2 represent an exponent or a multiplier or something different? If it's an exponent, write ^2 to indicate exponentiation. @pgpilot326 : Have to ask you to defend your " |x|^2 = a^2 + b^2 " ... where did that come from? |x| refers to "the absolute value of x."

OpenStudy (anonymous):

its not a exponent

OpenStudy (anonymous):

no 13 - 9 = ? that's what b^2 is

OpenStudy (anonymous):

you sure?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\left| x \right|^2\]right?

OpenStudy (anonymous):

no

OpenStudy (mathmale):

@elocke: please explain what " |x|2 = 13 " means to you. Use the correct mathematical vocabulary.

OpenStudy (anonymous):

\[\left| x \right|2\]?

OpenStudy (anonymous):

o wait it is

OpenStudy (anonymous):

sorry

OpenStudy (mathmale):

@pgpilot: sorry, but |x|2 isn't any clearer. What math operation(s) is or are involved here?

OpenStudy (mathmale):

|x| is "the absolute value of x." What is the purpose / function of that '2' ?

OpenStudy (mathmale):

@elocke: we're not getting anywhere. Please, would you DRAW your homework problem, using the Draw utility, below? The confusion here comes from your typing " |x|2 = 13 " ... which is not correct mathematically.

OpenStudy (mathmale):

For example, do you mean: \[|x|^{2}\]

OpenStudy (mathmale):

??

OpenStudy (mathmale):

sorry, I have to get off the 'Net, but I'll be back on later. Good luck.